#2: (2 pts.) Complete the following FULILAI IDDIUNT UYUL / 0 235 92 144 Xe 4 Sr + 54 38 1. ( inne into Au (s) in
Answer #1

+ 3 140 Bat 56 on - 2350 + In o 92 mass no = (235+1)-( 140+ x + 3) = 236 - (140tn-3) = 206 93 x Atomic no. = (92+0)-( 56tato)